OpenStudy (meilendurcer):

1. Evaluate: [(51 + 3) - 3^2] ÷ 9 • 2 2. Simplify the following expression. 2x - 6y + 3x^2 + 7y - 14x.

OpenStudy (mathstudent55):

Start working on the inner most parentheses first. Remember to copy everything else down so you don't forget part of the problem.

OpenStudy (meilendurcer):

@mathstudent55 I am :)

OpenStudy (mathstudent55):

Great, you're doing well.

OpenStudy (meilendurcer):

is it 10?

OpenStudy (calculusxy):

[(51 + 3) - 3^2] ÷ 9 • 2= =[(54)-9]/9*2 =45/9*2 =5*2 =10

OpenStudy (calculusxy):

So you're correct!

OpenStudy (meilendurcer):

Yay! ok can you help with 2?

OpenStudy (calculusxy):

Combine like terms to get: 3x2−12x+y

OpenStudy (meilendurcer):

@calculusxy ok im doing that

OpenStudy (calculusxy):

I think that this is the answer so you don't need to do anything else.